[figure: 9.7 cm]

OBSERVATION OF THE CHARGE-ORDERED STATE IN A

HETEROGENEOUS La_{0.7}Ca_{0.3}MnO_{3−δ} THIN FILM WITH OXYGEN DEFICIENCY

V. G. Prokhorov^{a}, Y. P. Lee^{b}, V. A. Komashko^{a}, V. Svechnikov^{a}, J. S. Park^{b}, and K. W. Kim^{c} aInstitute of Metal Physics, NASU, Kiev, 03142, Ukraine bDepartment of Physics, Hanyang University, Seoul, 133-791 Korea cDepartment of Physics, Sunmoon University, Choongnam, 336-840 Korea

The magnetic and transport investigations were carried out for the oxygen-deficit La_{0.7}Ca_{0.3}MnO_{3−δ} films (δ ≃ 0.1) with different thickness. The x-ray diffraction and high-resolution electron microscopy data showed that a growth in the film thickness leads to a structural transition from the lattice-strained cubic (with out-of-plane lattice parameter c ≃ 0.379 nm) to the relaxed rhombohedral phase (c ≃ 0.395 nm). The evidence for the charge-ordered insulating (COI) state was observed for the compressed cubic phase which is formed in the thin (d ≃ 30 nm) film owing to the epitaxial growth and to a large mismatch in the lattice parameter between substrate (LaAlO_{3}) and film. The applied magnetic field (5T) melts the COI state and leads to the transition of thin film in the ferromagnetic metal-like state. The magnetic data showed that the Curie point is shifted from ≃ 80 K for rhombohedral phase to ≃ 210 K for cubic one. It was claimed that formation of the COI state and shift of Curie point are governed by a compression of the crystal lattice rather than the accumulation of the oxygen vacancies.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : V. G. Prokhorov

Address for correspondence : Institute of Metal Physics, NASU, 36 Vernadsky Blvd., 03142 Kiev, Ukraine

Email address : pvg@imp.kiev.ua